AMINAT is 25yrs old in 1960 and her father is 35yrs older than her. How many years from now will the product of their ages be 2244
Answer by jorel1380(3719)

Let n be the number of years from 1960. Then:
(25+n)(60+n)=2244
nē+85n-744=0
(n+93)(n-8)=0
n=-93 or 8
Ignoring the negative result, we get a value of 8 years from 1960, or 1968 as the year the product of their ages is 2244. ☺☺☺☺